A cylinder contains a hydrogen gas of volume 2.40  103 m3 at 17 C and 2.32  106
Pa. Calculate
a. the number of molecules of hydrogen in the cylinder,
b. the mass of the hydrogen,
c. the density of hydrogen under these conditions.
(Given the molar mass of hydrogen = 2 g mol1
)
1
Expert's answer
2021-08-26T19:04:12-0400

a)


"N=\frac{(2.4\cdot10^{-3})(2.32\cdot10^{6})}{(1.38\cdot10^{-23})(273+17)}=1.39\cdot10^{24}"

B)


"m=\frac{(2.4\cdot10^{-3})(2.32\cdot10^{6})}{(8.31)(273+17)}2=4.62\ g"

C)


"\rho=\frac{4.62\cdot10^{-3}}{2.4\cdot10^{-3}}=1.93\frac{kg}{m^3}"